# Large Language Model for Vulnerability Detection and Repair: Literature Review and Roadmap

### v1

@@ 7. The Road Ahead
In this section, we discuss current challenges (Section 7.1) and highlight promising opportunities (Section 7.2) for applying LLMs techniques in vulnerability detection and repair.
7.1. Challenges
Challenge 1: Lack of High-quality Vulnerability Dataset. One major challenge is the lack of high-quality vulnerability benchmarks. Previous studies (Croft et al., 2023; Nie et al., 2023; Chen et al., 2023a) have highlighted issues with existing vulnerability data, including noisy or incorrect labels (e.g., labeling clean code as vulnerable). This data quality issue is primarily attributed to the use of automatic vulnerability collection (Nie et al., 2023), which can gather large enough data for training DL-based models including LLMs but cannot ensure the complete correctness of the labels. While manually checking each data sample can ensure high quality, it is a very tedious and expensive process, especially when aiming for a large dataset. Constructing a high-quality vulnerability benchmark remains an open challenge to date. Moreover, there is a growing concern that the capabilities of LLMs may derive from the inclusion of evaluation datasets in the pre-training corpus of LLMs, a phenomenon known as data contamination (Jiang et al., 2024). To mitigate this concern, high-quality vulnerability benchmarks are preferred to have no overlap with the pre-training corpus of LLMs.
Challenge 2: Complexity in Vulnerability Data. Vulnerabilities can be inherently complicated, which brings challenges for their detection and repair with LLMs. For instance, inter-procedural vulnerabilities are prevalent in vulnerability data, and Li et al. (Li et al., 2024) discovered that detecting inter-procedural vulnerabilities poses greater challenges than intra-procedural ones. Moreover, vulnerabilities encompass a wide array of Common Weakness Enumeration (CWE) types (Zhou et al., 2023a), but LLMs may struggle with less frequent CWE types compared to frequent types (Zhou et al., 2024a, 2023a). In addition, vulnerabilities are typically represented in terms of code units, such as code lines, functions, or program slices within which the vulnerabilities occur. Sejfia et al. (Sejfia et al., 2024) observed a significant accuracy drop when detecting vulnerabilities that span multiple code units, such as spanning multiple functions. Future research should consider the complex nature of vulnerabilities when designing LLM-based solutions.
@@ 7.1. Challenges
Challenge 3: Narrow Scope of Input Programs. Currently, LLM-based vulnerability detection and repair solutions primarily target the function level (or more fine-grained, at the line level). However, this narrow focus indicates that these approaches may not perform optimally when presented with a wider range of programs, such as classes or a whole repository. Function-level vulnerability detection approaches could overlook vulnerabilities that span multiple functions or classes (Sejfia et al., 2024). Similarly, function-level vulnerability repair approaches fall short when tasked with modifications across multiple functions within the repository (Zhou et al., 2024a). In addition to functions, future research could propose LLM-based detection/repair approaches capable of handling a broader range of programs.
@@
Challenge 4: High Accuracy and Robustness. A vulnerability detection or repair solution with high accuracy is generally preferred, as it boosts developers’ confidence in the reliability of detections and fixes. However, current state-of-the-art approaches (Liu et al., 2024; Zhou et al., 2024a) have not yet achieved satisfactory accuracy, with 67.6% and 20% accuracy scores for vulnerability detection and repair, respectively. Moreover, the solution should maintain robustness against data perturbations or adversarial attacks to ensure its resilience. However, Yang et al. (Yang et al., 2022) and Rahman et al. (Mahbubur Rahman et al., 2024) discovered that LLMs are not robust against data perturbations. Future research should seek ways to improve the accuracy and robustness of LLM-based solutions.
Challenge 5: Establishing Trust and Synergy with Developers. To enhance developers’ trust, LLM-based approaches should explicitly notify developers when unable to provide accurate outputs, which allows developers to avoid spending time reviewing potentially inaccurate outputs (Le et al., 2015). In addition, there is limited interaction between developers and LLM-based vulnerability detection/repair solutions, which may hinder the establishment of trust and synergy during practical application. To bridge this gap, future research should investigate more effective strategies for fostering collaboration and trust between developers and LLM-based solutions (Lo, 2023). By nurturing trust and synergy, LLM-based solutions have the potential to evolve into intelligent allies, offering enhanced support to developers.
7.2. Opportunities
@@
Opportunity 1: Curating High-quality Test Set. The absence of a high-quality vulnerability dataset poses a significant obstacle to vulnerability detection. While obtaining fully correct labels for a large dataset is expensive, a viable solution is to curate a high-quality test set (which is much smaller than the whole dataset) that can accurately assess progress in vulnerability detection. An easy approach for future work is to combine the manually checked vulnerability data samples scattered across several separate research works (e.g., (Croft et al., 2023; He and Vechev, 2023; Li et al., 2023b)) to form a high-quality test set. Future works can then do an empirical study to recognize the real progress in vulnerability detection with the high-quality test set. Furthermore, the community can maintain a living high-quality test set by adding new manually verified data to it. This curated test set can serve as a reliable benchmark for vulnerability detection.
@@
Opportunity 2: Repo-level Vulnerability Detection/Repair. Current vulnerability detection and repair techniques primarily focus on the function or line level. One key reason is the input length limitations (512 subtokens) of the small LLMs like CodeBERT and CodeT5, which are predominantly used in existing studies. The 512 subtokens limitation aligns well with the function level data but faces difficulty in scaling up to classes or repositories. However, the advent of recent larger LLMs with significantly higher input length capacity, such as Code Llama (16,384 subtokens) and GPT-4 (24,576 sub-tokens), enables processing repo-level data more effectively. This presents an opportunity for future research to explore repo-level vulnerability detection/repair by leveraging these larger LLMs.
Opportunity 3: Using Larger Decoder-only LLMs. As presented in RQ1, it is evident that researchers favor encoder-only language models (61.5%) for vulnerability detection and encoder-decoder models (48.0%) for repair. Moreover, only 6 out of the 36 studies investigate LLMs with more than 7 billion parameters, suggesting a focus on smaller LLMs. This preference contrasts with a broader trend observed in LLMs for SE, as highlighted in a recent literature review (Hou et al., 2023). They found that since 2023, decoder-only models have become predominant, accounting for 73.1% of the papers studied in 2023. Additionally, among very large LLMs (e.g., ¿7B parameters), the majority are decoder-only models. Furthermore, decoder-only LLMs excel in generating code/text (OpenAI, 2023), making them well-suited for vulnerability repair, which is a generation task. While vulnerability detection is a classification task, it can be reframed as a generation task. For instance, decoder-only LLMs can be instructed to output ‘it is vulnerable’ given a vulnerable code and ‘it is non-vulnerable’ given a non-vulnerable code. Thus, large decoder-only LLMs can be leveraged to address both vulnerability detection and repair tasks. Currently, it seems that decoder-only LLMs, particularly those with parameters larger than 7 billion, are underutilized in vulnerability detection and repair, indicating untapped potential and offering opportunities for future work.
Opportunity 4: Advanced LLM Usage and Adaptation. Regarding LLMs usages, beyond the techniques observed in the included studies—such as fine-tuning and zero/few-shot prompting—there exist two more advanced usages of LLMs (Minaee et al., 2024) that have not been explored yet in vulnerability detection and repair: 1) LLM Agent: LLMs can serve as agents to decompose complex tasks into smaller components and employ multiple LLMs to address them (Wang et al., 2024a); 2) Usage of External Tools: LLMs can utilize external tools such as search engines, external databases, and other resources to enhance them (Qin et al., 2023).
Regarding detailed LLMs adaptation techniques, as illustrated in Fig.5 and Fig.7, the majority of proposed adaptations in this field are designed for fine-tuning. However, a plethora of advanced adaptation techniques for few-shot prompting remains unexplored. These techniques include iterative retrieval augmentation (Shao et al., 2023), recursive retrieval augmentation (Trivedi et al., 2022), and adaptive retrieval augmentation (Asai et al., 2023). Researchers can consider using those unexplored advanced LLM usages/adaptations in future works.
@@
Opportunity 5: Customized LLMs for Vulnerability. Currently, widely used LLMs in vulnerability detection/repair are general-purpose LLMs (e.g., CodeBERT, CodeT5, and GPT-3.5) that do not fully exploit the wealth of open-source vulnerability data. A promising avenue is the development of customized LLMs tailored for vulnerability data. Some initial attempts in this direction include vulnGPT (Vicarius, 2023) and Microsoft Security Copilot (Microsoft, 2024). However, as these solutions are proprietary, their customized LLM details may not be fully disclosed. We advocate for collaborative efforts to develop open-sourced and effective customized LLMs for vulnerability.
@@ 9. Conclusion and Future Work
The use of Large Language Models (LLMs) for vulnerability detection and repair has been garnering increasing attention. This paper presents a systematic literature review of 36 primary studies on LLMs for vulnerability detection and repair, sourced from leading SE, AI, and Security conferences and journals. This review begins by analyzing the types of LLMs used in primary studies, shedding light on researchers’ preferences for different LLMs. Subsequently, we categorized a variety of techniques for adapting LLMs. Through our analysis, this review also identifies the challenges in this field and proposes a research roadmap outlining promising avenues for future exploration. In the future, we plan to broaden this literature review by incorporating additional vulnerability-related tasks, such as vulnerability localization and vulnerability assessment. Our online appendix is available at (app, 2024).

### v3

@@ 8. The Road Ahead
In this section, we first discuss the limitations of current studies and then introduce a roadmap, illustrating the research trajectory shaped by prior work and highlighting future directions for exploration.
8.1. Limitations
@@
Limitation 1: Small Input Granularity. Currently, LLM-based vulnerability detection and repair solutions primarily target the function level (or more fine-grained, at the line level). However, this small input granularity indicates that these approaches may not perform optimally when presented with a wider range of programs, such as classes or a whole repository. Function-level vulnerability detection approaches could overlook vulnerabilities that span multiple functions or classes (sejfia2024toward, ). Similarly, function-level vulnerability repair approaches fall short when tasked with modifications across multiple functions within the repository (zhou2024out, ). In addition to functions, future research could propose LLM-based detection/repair approaches capable of handling a broader range of programs.
Limitation 2: Lack of High-quality Vulnerability Dataset. One major challenge is the lack of high-quality vulnerability datasets. For vulnerability repair, previous studies (DBLP:conf/icse/CroftBK23, ; DBLP:conf/issta/NieLWWLW23, ; DBLP:conf/raid/0001DACW23, ) have highlighted issues with existing vulnerability data, including noisy or incorrect labels (e.g., labeling clean code as vulnerable). This data quality issue is primarily attributed to the use of automatic vulnerability collection (DBLP:conf/issta/NieLWWLW23, ), which can gather large enough data for training DL-based models including LLMs but cannot ensure the complete correctness of the labels. While manually checking each data sample can ensure high quality, it is a very tedious and expensive process, especially when aiming for a large dataset. Constructing a high-quality vulnerability detection benchmark remains an open challenge to date.
For vulnerability repair, most studies either utilize synthetic data (e.g.,  (islam2024code, )) or use real-world data without test cases (e.g.,  (DBLP:journals/tse/ChenKM23, ; DBLP:journals/tse/ChiQLZY23, ; DBLP:conf/sigsoft/FuTLNP22, )). The reliance on synthetic data can limit the generalizability of the findings, as it may not accurately reflect the complexities and nuances of actual vulnerabilities encountered in production environments. On the other hand, real-world data without test cases makes it difficult to evaluate the effectiveness of proposed solutions comprehensively, as the absence of test cases hampers the ability to verify the correctness of the vulnerability repairs. There is an urgent need for large real-world vulnerability-fixing datasets that include test cases.
Moreover, there is a growing concern that the capabilities of LLMs may derive from the inclusion of evaluation datasets in the pre-training corpus of LLMs, a phenomenon known as data contamination (jiang2024investigating, ). To mitigate this concern, high-quality vulnerability detection/fixing benchmarks are preferred to have no overlap with the pre-training corpus of LLMs.
Limitation 3: Suboptimal Performance Caused by Complexity in Vulnerability Data. Vulnerabilities can be inherently complicated, which brings challenges for their detection and repair with LLMs. For instance, inter-procedural vulnerabilities are prevalent in vulnerability data, and Li et al. (li2024effectiveness, ) discovered that detecting inter-procedural vulnerabilities poses greater challenges than intra-procedural ones. Moreover, vulnerabilities encompass a wide array of Common Weakness Enumeration (CWE) types (zhou2023devil, ), but LLMs may struggle with less frequent CWE types compared to frequent types (zhou2024out, ; zhou2023devil, ). In addition, vulnerabilities are typically represented in terms of code units, such as code lines, functions, or program slices within which the vulnerabilities occur. Sejfia et al. (sejfia2024toward, ) observed a significant accuracy drop when detecting vulnerabilities that span multiple code units, such as spanning multiple functions. Future research should consider the complex nature of vulnerabilities when designing LLM-based solutions.
Limitation 4: Reliance on Lightweight LLMs. As highlighted in Table 2 and Table 3, most studies adapted lightweight LLMs (#parameter ¡ 1B) for vulnerability detection and repair. For lightweight LLMs like CodeBERT, researchers have explored various strategies to boost their performance, including data-centric enhancements, model-centric innovations, integration with program analysis, combining LLMs with other deep learning methods, domain-specific pre-training, causal learning, and reinforcement learning. In some cases, such as Liu et al. (liu2024pre, ), multiple techniques have been combined to achieve even greater performance, highlighting the versatility and potential of lightweight models. In contrast, the use of large LLMs (with over 1 billion parameters) remains relatively limited compared to lightweight LLMs, leaving a large room to explore.
Limitation 5: Lack of Deployment Consideration. We examine two crucial aspects of deployment: 1) interaction with developers and 2) integration into developers’ current workflows. Currently, none of the studies in our review have been incorporated into developers’ workflows; instead, they primarily rely on static and historical data to evaluate their effectiveness. Moreover, limited studies in our review incorporate interaction with developers, such as engaging with developers through feedback or explaining the rationale behind vulnerability detection. Limited interaction between developers and LLM-based solutions may impede the establishment of trust and collaboration during practical applications. This lack of interactive features restricts the practical applicability of these models in real-world scenarios, where effective vulnerability management relies on collaboration between automated tools and developers.
To address these gaps, future research should explore more effective strategies for fostering collaboration and trust between developers and LLM-based solutions (DBLP:journals/corr/abs-2309-04142, ). Additionally, seamlessly integrating LLM-based solutions can enable them to evolve into intelligent partners, providing enhanced support to developers.
@@
Limitation 6: Lack of High Accuracy and Robustness. A vulnerability detection or repair solution with high accuracy is generally preferred, as it boosts developers’ confidence in the reliability of detections and fixes. However, current state-of-the-art approaches (liu2024pre, ; zhou2024out, ) have not yet achieved satisfactory accuracy, with 67.6% and 20% accuracy scores for vulnerability detection and repair, respectively. Moreover, the solution should maintain robustness against data perturbations or adversarial attacks to ensure its resilience. However, Yang et al. (yang2022natural, ) and Rahman et al. (mahbubur2023towards, ) discovered that LLMs are not robust against data perturbations. Future research should seek ways to improve the accuracy and robustness of LLM-based solutions.
Figure 8. Roadmap and Future Directions
8.2. Roadmap
Figure 8 presents the roadmap. The green modules highlight the aspects that existing studies have primarily focused on, while the gray modules represent areas that have not been extensively explored but warrant further investigation. The solid black lines represent the trajectories followed by most existing LLM-based methods for vulnerability detection and repair, while the gray dotted lines highlight relatively under-explored paths that could offer promising directions for future research in vulnerability detection and repair.
8.2.1. Established Pathways in Current Studies.
As shown in both Table 23 and the solid black lines in Figure 8, the existing methods typically leverage either lightweight LLMs or large LLMs adapted using techniques such as fine-tuning or prompting. The existing methods often focus on method-level or line-level analysis. The datasets used are usually collected from heuristic approaches for vulnerability detection or real-world data without test cases for vulnerability repair. Finally, these methods usually neglect deployment considerations, such as ensuring effective interaction with developers.
8.2.2. Path Forward.
As shown in Figure 8, there are many promising yet under-explored modules (highlighted in green) and paths (represented by the gray dotted lines.) Looking ahead, we can outline three progressively advanced stages for exploring new directions. In the first stage, researchers can study these under-explored modules individually. In the second stage, they can investigate the under-explored paths, which may include combinations of under-explored modules and well-established ones or consist solely of multiple under-explored modules. Finally, in the third stage, by transforming all under-explored modules into well-understood components and studying various under-explored paths and combinations, researchers can summarize their findings and analyze strategies for better developing LLM-based solutions, ultimately leading to the new generation of powerful LLM-based vulnerability detection and repair solutions. Below, we describe the three stages in detail.
Stage 1: Exploring Under-explored Modules Individually. Figure 8 highlights several under-explored modules in green, each offering exciting potential for discovery. We outline below some promising opportunities for deeper investigation into these areas:
@@ 8.2. Roadmap
  • Opportunity 1: Curating High-quality Test Set for Vulnerability Detection. The absence of a high-quality vulnerability dataset poses a significant obstacle to vulnerability detection. While obtaining fully correct labels for a large dataset is expensive, a viable solution is to curate a high-quality test set (which is much smaller than the whole dataset) that can accurately assess progress in vulnerability detection. An easy approach for future work is to combine the manually checked vulnerability data samples scattered across several separate research works (e.g., (DBLP:conf/icse/CroftBK23, ; he2023large, ; li2023comparison, )) to form a high-quality test set. Future works can then do an empirical study to recognize the real progress in vulnerability detection with the high-quality test set. Furthermore, the community can maintain a living high-quality test set by adding new manually verified data to it. This curated test set can serve as a reliable benchmark for vulnerability detection.
@@
  • Opportunity 2: Repo-level Vulnerability Detection/Repair. Current vulnerability detection and repair techniques primarily focus on the function or line level. One key reason is the input length limitations (512 subtokens) of the small LLMs like CodeBERT and CodeT5, which are predominantly used in existing studies. The 512 subtokens limitation aligns well with the function level data but faces difficulty in scaling up to classes or repositories. However, the emergence of recent larger LLMs with significantly higher input length capacities, such as GPT-4 (which can handle 128k subtokens), facilitates more effective processing of repo-level data. This presents an opportunity for future research to explore repo-level vulnerability detection/repair by leveraging these larger LLMs. For this direction, one recent study (zhou2024comparison, ) has initially explored the repo-level vulnerability detection task. Specifically, it compared SAST tools (e.g., CodeQL) with popular or state-of-the-art open-source LLMs (e.g., Code Llama) for detecting software vulnerabilities in software repositories. The experimental results indicated that SAST tools achieve low vulnerability detection rates with relatively low false positives, whereas LLMs can detect more vulnerabilities but tend to suffer from high false positive rates. Alongside this initial effort, there is considerable potential and opportunities for further research in this direction.
@@
  • Opportunity 3: Customized LLMs for Vulnerability. Currently, widely used LLMs in vulnerability detection/repair are general-purpose LLMs (e.g., CodeBERT, CodeT5, and GPT-3.5) that do not fully exploit the wealth of open-source vulnerability data. A promising avenue is the development of customized LLMs tailored for vulnerability data. Some initial attempts in this direction include vulnGPT (vul_GPT, ) and Microsoft Security Copilot (SecurityCopilot, ). However, as these solutions are proprietary, their customized LLM details may not be fully disclosed. We advocate for collaborative efforts to develop open-sourced and effective customized LLMs for vulnerability.
  • Opportunity 4: Advanced LLM Usage and Adaptation. Regarding LLMs usages, beyond the techniques observed in the included studies—such as fine-tuning, prompt engineering, and retrieval augmentation—there exist two more advanced usages of LLMs (minaee2024large, ) that have not been explored yet in vulnerability detection and repair: 1) LLM Agent: LLMs can serve as agents to decompose complex tasks into smaller components and employ multiple LLMs to address them (wang2024survey, ); 2) Usage of External Tools: LLMs can utilize external tools such as search engines, external databases, and other resources to enhance them (qin2023toolllm, ). Regarding detailed LLMs adaptation techniques, as illustrated in Fig.6 and Fig.7, the majority of proposed adaptations in this field are designed for fine-tuning and prompt engineering. However, a plethora of advanced adaptation techniques for retrieval augmentation remains under-explored. These techniques include iterative retrieval augmentation (shao2023enhancing, ) and recursive retrieval augmentation (trivedi2022interleaving, ), and adaptive retrieval augmentation (asai2023self, ). Researchers can consider using those unexplored advanced LLM usages/adaptations in future works.
  • Opportunity 5: Support Deployment-ready Features. Efforts could focus on enhancing LLM-based solutions with deployment-ready features, such as user interaction capabilities and seamless integration into existing developer workflows. This includes developing intuitive interfaces that allow developers to provide real-time feedback on suggestions made by the LLMs, thereby creating a more collaborative environment. Additionally, implementing functionalities that explain the reasoning behind the model’s recommendations can help demystify the decision-making process, fostering trust and facilitating better adoption among developers. Moreover, integrating these solutions into popular IDEs or version control systems can streamline the workflow, making it easier for developers to utilize LLMs as part of their daily practices. By prioritizing these enhancements, LLM-based tools can significantly improve their usability and effectiveness in real-world scenarios, ultimately contributing to more robust vulnerability detection and repair processes.
Stage 2: Exploring Under-explored Paths. The under-explored paths (mainly represented by gray dotted lines from data to deployment) may include combinations of under-explored modules with well-established ones or consist solely of multiple under-explored modules. This stage presents numerous opportunities. Below, we first highlight some examples of studying these combinations of under-explored modules alongside well-established ones:
  • Once the community develops customized LLMs, researchers can apply well-studied fine-tuning or prompting-based adaptation techniques to improve function-level or line-level vulnerability detection or repair. These enhanced models can further be extended to support developer interaction, creating more practical and integrated solutions.
  • Once new datasets are established—such as those containing manually verified vulnerabilities and associated test cases—researchers can re-evaluate existing LLM methods to assess their effectiveness in vulnerability detection and repair. Exploring various combinations of under-explored aspects and well-studied approaches can yield valuable insights and advancements.
  • Researchers can leverage general-purpose LLMs (e.g., CodeBERT, CodeT5, and GPT-3.5) to develop repo-level vulnerability detection and repair tools, also facilitating interaction with developers.
There are also many opportunities in paths that consist solely of multiple under-explored modules. For instance, researchers could design a repo-level vulnerability detection or repair method using customized vulnerability-specific LLMs and enhance its effectiveness by incorporating LLM agent techniques. Finally, the tool can be integrated into IDEs, leveraging LLMs to improve interaction capabilities.
Stage 3: Summarizing Findings and Applying Them for Next-Generation LLM-based Solutions. In this most advanced stage, researchers are expected to transform various under-explored aspects into well-studied areas and investigate diverse under-explored paths. At this stage, they can summarize their findings and analyze strategies for improving the development of LLM-based solutions. By leveraging this experience and knowledge, they may push the boundaries of what LLMs can achieve in vulnerability detection and repair to new heights.
Ideally, by 2030, following the roadmap illustrated in Figure 8, the community can develop a highly effective method capable of performing line-level, method-level, class-level, and repository-level vulnerability detection and repair, utilizing an expert-annotated benchmark to ensure accurate evaluation. Furthermore, the ideal solution should facilitate seamless communication and collaboration with developers, achieving high accuracy and robustness while providing trustworthy, expert-level insights into real-world vulnerabilities.
@@ 10. Conclusion and Future Work
The use of Large Language Models (LLMs) for vulnerability detection and repair has been garnering increasing attention. This paper presents a systematic literature review of 58 primary studies on LLMs for vulnerability detection and repair. This review begins by analyzing the types of LLMs used in primary studies, shedding light on researchers’ preferences for different LLMs. Subsequently, we categorized a variety of techniques for adapting LLMs. Through our analysis, this review also identifies the limitations in this field and proposes a research roadmap outlining promising avenues for future exploration. In the future, we plan to broaden this literature review by incorporating additional vulnerability-related tasks, such as vulnerability localization and vulnerability assessment.
Acknowledgement. This research / project is supported by the National Research Foundation, under its Investigatorship Grant (NRF-NRFI08-2022-0002). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.